OpenStudy (anonymous):

Shonda made a model of a solid ice cream cone with a spherical scoop of ice cream on top, made out of clay. Which formulas are needed to calculate how much clay she used? Choose all answers that are correct. (In these formulas, r represents radius, h represents height, l represents length, and w represents width.) A. B. C. D.

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

you may have noticed that nothing posted in your post ....

OpenStudy (amistre64):

but if we have a cone and sphere ... then the volume formula for them will be useful to us

OpenStudy (anonymous):

A. how do i attch the whole file

OpenStudy (amistre64):

take a picture of it and post the picture file, screenshot or something

OpenStudy (amistre64):

or just type them in

OpenStudy (anonymous):

Shonda made a model of a solid ice cream cone with a spherical scoop of ice cream on top, made out of clay. Which formulas are needed to calculate how much clay she used. Choose all answers that are correct. A. v =1/3 pi r 2/n B. v =pi r 2/h C. v = Iwh D. v = 4/3 pi r 3 ( in these formulas, r represents radius, h represents height, and l represents length and the w represents the width.)

OpenStudy (amistre64):

thats better now, what do you recall of the volume formulas for a sphere and a cone?

OpenStudy (anonymous):

A and D

OpenStudy (amistre64):

good, and C is a box, B is a cylindar which is technically part of A B may or maynot be an answer choice depending on how they are grading it

OpenStudy (anonymous):

so what you mean technically part of A

OpenStudy (amistre64):

the volume of a cone is 1/3 the volume of its representative cylindar. hence: cylindar = pi r^2 h cone = 1/3 cylindar = 1/3 pi r^2 h

OpenStudy (anonymous):

so it is a and d right

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

a and d for sure, but the way the question is worded, b may or maynot be included ... i cant be certain of it

OpenStudy (anonymous):

Two spherical balloons are filled with water. The first balloon has a radius of 3 cm, and the second has a radius of 6 cm. How much more water is in the larger balloon than in the smaller balloon? Use 3.14 to approximate pi, and express your answer in the hundredths.

OpenStudy (anonymous):

can you help me

OpenStudy (amistre64):

use the volume formula for the sphere ... wht do we get?

OpenStudy (anonymous):

which one do i do first 3 or 6 radius

OpenStudy (amistre64):

both of them, since we need to know the difference between them

OpenStudy (amistre64):

or we could try to work out a general formula like: V1 = 4/3 pi r1^3 V2 = 4/3 pi r2^3 V1 - V2 = 4/3 pi (r1^2-r2^3)

OpenStudy (anonymous):

i thought they were both cubed

OpenStudy (amistre64):

yeah, slip of the fingers :)

OpenStudy (anonymous):

so it is 791.28

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

4/3 (3.14) (6^3-3^3) http://www.wolframalpha.com/input/?i=4%2F3+%283.14%29+%286%5E3-3%5E3%29 yep

OpenStudy (amistre64):

gotta brb